OpenStudy (lgbasallote):

LGBADERIVATIVE: \[F(t) = e^{t\sin 2t}\]

OpenStudy (lgbasallote):

im thinking implicit differentiation?

OpenStudy (lgbasallote):

\[\huge \ln y = t\sin 2t\]

OpenStudy (lgbasallote):

\[\frac 1y y' = 2t\cos (2t) + \sin(2t)\]

OpenStudy (anonymous):

Chain Rule, that's all you need

OpenStudy (lgbasallote):

\[y' = e^{t\sin 2t} (2t\cos (2t) + \sin (2t) )\]

OpenStudy (lgbasallote):

is that right?

OpenStudy (anonymous):

yes.

OpenStudy (apoorvk):

yes that is correct.

OpenStudy (lgbasallote):

great thanks :D

OpenStudy (anonymous):

is that a new term? "LGBADERIVATIVE" i can't find it in my dictionary...:(

OpenStudy (lgbasallote):

nowadays...everything is named after lgba :p

OpenStudy (anonymous):

ah..... :)